Please help could I get an explanation of the answer?
7(2x - 3)=20x

Answer by stanbon(75887)

7(2x - 3)=20x
Use distributive property on the left side.
7*2x - 7*3 = 20x
Simplify the left side.
14x - 21 = 20x
Subtract 14x from both sides to get:
6x = -21
Divide both sides by 6 to get:
x = -7/2
=====================
Cheers,
Stan H.